OpenStudy (anonymous):

FIND 3RD DERIVATE (X) / (X+1)

OpenStudy (anonymous):

y' = [(x + 1)(1) - (x)(1)]/(x + 1)^2 = 1/(x + 1)^2 y'' = -2(x + 1)^-3 y''' = 6(x + 1)^-4 = 6/(x + 1)^4

OpenStudy (anonymous):

;=)